OpenStudy (anonymous):

Find the smallest number that is divisible by both 3 and 7 and also the sum of the digits of this number is divisible by 3 and 7. For example, 42 is divisible by 3 and 7 but 4+2=6 is not divisible by 7. Therefore, 42 is not the answer. The answer is a 3-digit number less than 500.

OpenStudy (amistre64):

well, it has to be a multiple of 21 for sure

OpenStudy (amistre64):

21^2 by chance?

OpenStudy (amistre64):

nah, 441 is 9

OpenStudy (amistre64):

21x = 100a+10b+c a+b+c = 21y just another random thought

OpenStudy (amistre64):

arent all 3 digit numbers less than 5000?

OpenStudy (anonymous):

sorry, its 500, not 5000, dumb computer XD thank you for helping :) this is for my shsat help so i can get into a good highschool

OpenStudy (amistre64):

none of the numbers are going to add to anyting much greater than 21 499 =22 at best

myininaya (myininaya):

so we have the number is a 3 digit number so we know we have p=21*k for some integer k>0 (i'm assuming smallest positive number of course) we also have p=d_2d_1d_0 so we have \[d_2 10^2+d_110^1+d_010^0=21 \cdot k \] we also have \[d_210^2+d_110^1+d_010^0=3 k_1 \] so we know \[d_1+d_2+d_3=3 \cdot k_2\] I don't know if there is a trick for 7. Also I put different k's above because they aren't equal k's

myininaya (myininaya):

actually you don't have to go far in the 100's to find the answer

myininaya (myininaya):

actually the fist number that takes in the 3 digits numbers is 21*5 and guess what?

myininaya (myininaya):

oops sum of digits sorry

myininaya (myininaya):

105 is divisible for 3 and 7 i forgot it said digits of the number

OpenStudy (amistre64):

21x = 100a+10b+c 21 = a+b+c 21 = (100a+10b+c)/x a+b+c = (100a+10b+c)/x ax+bx+cx = 100a+10b+c 0 = (100-x)a+(10-x)b+(1-x)c maybe we can go with another thought ............. 21x = 100a+10b+c 21 = a+b+c x = 100a + 10b + c -------------- a+b+c the only way i can think of is to make sure is to brute math it lol

OpenStudy (amistre64):

there are only 21 numbers to deal with that are less than 500

OpenStudy (amistre64):

21=3 42=6 63=9 84=12 105=6 126=9 147=12 168=15 189=18 210=3 231=6 252=9 273=12 294=15 315=9 336=12 357=15 378=18 399=21 ..... this is it 420 441 462 483

OpenStudy (amistre64):

fine 23 numbers lol

ganeshie8 (ganeshie8):

yeah subtracting both equations eliminates c and the problem simplifies to : 100a+10b+c = 21k a+b+c = 21y ---------------------- 99a+9b = 21m

OpenStudy (amistre64):

y = 1